### v2

# RxnBench: A Multimodal Benchmark for Evaluating Large Language Models on Chemical Reaction Understanding from Scientific Literature

Abstract
The integration of Multimodal Large Language Models (MLLMs) into chemistry promises to revolutionize scientific discovery, yet their ability to comprehend the dense, graphical language of reactions within authentic literature remains underexplored. Here, we introduce RxnBench, a multi-tiered benchmark designed to rigorously evaluate MLLMs on chemical reaction understanding from scientific PDFs. RxnBench comprises two tasks: Single-Figure QA (SF-QA), which tests fine-grained visual perception and mechanistic reasoning using 1,525 questions derived from 305 curated reaction schemes, and Full-Document QA (FD-QA), which challenges models to synthesize information from 108 articles, requiring cross-modal integration of text, schemes, and tables. Our evaluation of MLLMs reveals a critical capability gap: while models excel at extracting explicit text, they struggle with deep chemical logic and precise structural recognition. Notably, models with inference-time reasoning significantly outperform standard architectures, yet none achieve 50% accuracy on FD-QA. These findings underscore the urgent need for domain-specific visual encoders and stronger reasoning engines to advance autonomous AI chemists.
Introduction
Figure 1: The RxnBench framework for advancing chemical reaction understanding. (a) Challenges in reaction process understanding. Interpreting reaction schemes transcends simple object detection, requiring models to infer dynamic roles—reactants, intermediates, and products—within complex chemical transformations. (b) Two-tiered evaluation hierarchy. RxnBench simulates a chemist’s cognitive workflow through: (i) Single-Figure QA (SF-QA) for granular visual perception and local mechanistic reasoning; and (ii) Full-Document QA (FD-QA) for cross-modal synthesis and reasoning within unstructured PDF documents.
The past decade has witnessed a paradigm shift in scientific discovery driven by Artificial Intelligence (AI), particularly in chemistry, the central science, where data-driven approaches are reshaping traditional research workflows Wang2023. From early Quantitative Structure-Activity Relationship (QSAR) models Cherkasov2014 to modern deep learning-based molecular generation GomezBombarelli2018 and property prediction Yang2019, AI has demonstrated potential that often transcends human intuition. Notably, the breakthrough in protein structure prediction achieved by geometric deep learning models like AlphaFold Jumper2021 has significantly inspired the chemical community to explore AI applications in small molecule drug discovery Vamathevan2019, materials design Merchant2023, and complex organic retrosynthesis planning Segler2018; Coley2019.
In recent years, the emergence of Large Language Models (LLMs) has introduced a new variable to chemical research. Unlike specialized models focused on single tasks, LLMs exhibit surprising general reasoning capabilities and knowledge integration skills OpenAI2023. Through instruction tuning or agent-based frameworks (such as ChemCrow Bran2023), LLMs have been proven capable of performing literature retrieval, experimental planning, and even controlling automated laboratory equipment Boiko2023. However, most existing chemical large models rely heavily on linearized text representations (such as SMILES Weininger1988 or SELFIES Krenn2022) to process chemical information. Although the textual modality is a significant carrier of knowledge, the core knowledge in chemical literature is often highly condensed within the unstructured visual modality, limiting the complete acquisition and understanding of chemical knowledge by text-only LLMs.
There are unique challenges in chemical literature understanding. Achieving machine reading of chemical literature faces multiple challenges not encountered in general document understanding tasks: First, the chemical reaction scheme is a highly condensed domain-specific graphical language. Unlike natural images, a reaction scheme is a complex semantic network composed of molecular structures, arrow symbols, reaction conditions (temperature, solvents, catalysts, etc.), and text labels. As shown in Figure 1a, understanding a reaction scheme requires not only recognizing discrete molecular structures but also inferring their roles in a dynamic chemical process (reactants, products, intermediates, or transition states). This type of “process understanding” is far more complex than simple object detection. Second, the unstructured nature of PDF documents poses significant engineering and cognitive hurdles. Scientific literature is primarily disseminated in PDF format—a format designed for human visual consumption rather than machine parsing—often referred to as a “graveyard” for data Tkaczyk2015. In chemical papers, reaction scheme images are frequently spatially separated from their textual descriptions, requiring models to handle cross-page layout dependencies. Third, the strong association and complementarity of multimodal information (Strong Multimodality). Chemical knowledge rarely exists in a single modality in isolation. For instance, a code in the main text (e.g., “Compound 3a”) must be aligned with the structure in an illustration (e.g., “Scheme 1”), while specific yields or enantiomeric excess ($e$) values might be hidden in a supplementary table (e.g., “Table S2”). Models require the ability to perform joint reasoning across text, images, and tables Yin2024. Finally, the extreme requirement for precision. Chemistry is an exact science. In natural image captioning, misidentifying a “dog” as a “cat” is an error; however, in chemistry, a mere change in the direction of a chemical bond (wedge vs. dash) implies a reversal of stereochemistry, which in medicinal chemistry can mean the difference between a cure and a poison (e.g., the thalidomide tragedy). Existing multimodal models suffer from severe “hallucination” issues Nori2023, tending to generate structures that appear plausible but are chemically incorrect.
As for the existing benchmarks for chemistry, we categorize them into two types based on input modality: single-modality benchmarks, which present problems purely in textual form, and multimodality benchmarks, which incorporate both textual and visual inputs, such as molecular structures or crystallographic images. Representative single-modal benchmarks include ChemBench Mirza2024, ChemLLMBench Guo2023, MMLU Hendrycks2020, SciBench Wang2023SciBench, SciCode Tian2024, JEEBench Arora2023, and OlympicArena He2024. While these provide robust evaluation for textual reasoning, they inherently lack the ability to assess visual chemical literacy. Conversely, multimodal initiatives such as ScienceQA Saikh2022, OlympiadBench He2024, ChemTable Zhou2025, Mol-Puzzle Guo2024, and ScemQA Liang2024 have integrated visual data to assess broader scientific literacy. Yet, a critical gap remains: these benchmarks primarily target static concept recognition or elementary interpretation, often treating molecular graphs as isolated objects rather than components of a dynamic reaction process. They generally lack the granularity required to evaluate the precise structural and quantitative reasoning—such as evaluating reaction yields of different products, or verifying stereochemical fidelity—that is central to organic synthesis. Furthermore, none of these datasets challenge models to synthesize information from the full, unstructured context of reaction-containing PDF literatures, a capability that fundamentally distinguishes our work from existing concept-centric evaluations. This limitation motivates the development of RxnBench, which aims to address these gaps through broader coverage of reaction types and a distinct focus on reaction perception and full-document reasoning.
Our Contributions. To bridge this gap, we introduce RxnBench, a benchmark designed specifically to evaluate the chemical reaction literature understanding capabilities of Multimodal Large Language Models (MLLMs). The core design philosophy of RxnBench is to simulate the authentic cognitive process of a chemist reading literature, comprising tasks at two levels (Figure 1b):
  1. (1)
    Single-Figure QA (SF-QA): Focuses on the “perception” level. We have constructed a dataset containing a high diversity of reaction types, requiring models to precisely extract reactant structures (SMILES), reaction conditions, yields, and stereochemical information from reaction scheme images. This aims to evaluate the model's parsing precision of the chemical graphical language.
  2. (2)
    Full-Document QA (FD-QA): Focuses on the “reasoning” level. This is the first chemical benchmark requiring models to answer questions within the context of complete PDF literature. Models must integrate main text descriptions, reaction mechanism diagrams, and experimental data tables to perform cross-modal information retrieval and comprehensive reasoning.
RxnBench aims to provide the community with a standardized evaluation tool, not only revealing the capability boundaries of current State-of-the-Art (SOTA) scientific multimodal models but also pointing the way toward the evolution of more precise and reasoning-capable “AI Chemists”.
Construction of RxnBench
Figure 2: Construction methodologies and evaluation logic of RxnBench. (a) SF-QA construction pipeline. A systematic workflow integrating automated extraction, human-in-the-loop refinement, and adversarial editing to ensure high-quality, chemically rigorous question-answer pairs. (b) FD-QA evaluation hierarchy. A comprehensive framework simulating a chemist’s real-world workflow, requiring multi-page synthesis and cross-modal reasoning across text, schemes, and tables within complete scientific articles.
To ensure RxnBench serves as a rigorous and representative testbed for modern chemical reasoning, we implemented a meticulous data curation and annotation pipeline, prioritizing high-quality sources and domain-expert verification.
Data Curation and Source Selection.
We constructed our corpus from high-impact scientific articles published within the last five years. To guarantee the diversity and complexity of the reaction data, we selected Open Access (OA) papers from eight prestigious journals that define the frontier of chemical research. These include top-tier multidisciplinary journals and leading specialized journals in organic and catalytic chemistry (Nature Chemistry, Nature Communications, JACS Au, Nature Synthesis, Journal of the American Chemical Society, Angewandte Chemie International Edition, ACS Catalysis, and Organic Letters). Our selection process focused on articles centered around organic reaction methodology and total synthesis. We filtered candidates to ensure every selected document contained:
  1. (1)
    Rich Graphical Information: At least one comprehensive reaction scheme depicting reactants, reagents, conditions, and products with clear stereochemical notation (if have).
  2. (2)
    Complex Text-Image Interplay: Full-text descriptions that necessitate cross-referencing with the figures (e.g., discussion of reaction mechanisms or optimization tables).
  3. (3)
    Chemical Diversity: A broad coverage of reaction classes, ranging from classic cross-couplings to contemporary photoredox catalysis and C–H activation, ensuring the benchmark is not biased toward simple textbook reactions.
Given the subtle nuances of organic chemistry—where a single stereocenter inversion or a misinterpreted abbreviation can invalidate a reaction—crowdsourcing platforms were deemed unsuitable for this task. Instead, we recruited a team of domain experts to perform the annotation. The annotation team consisted of doctoral candidates (Ph.D. students) specializing in organic chemistry. All annotators possessed extensive experience in reading synthetic literature and were proficient in using chemical drawing software and interpreting mechanism of organic reactions. Furthermore, they were skilled in Optical Chemical Structure Recognition (OCSR) workflows, to convert graphical structures directly into Extended SMILES (E-SMILES) Fang2024, facilitating precise structural editing and verification.
Construction of Single-Figure QA (SF-QA)
The construction of the SF-QA dataset followed a systematic workflow designed to ensure both scalability and depth (Figure 2a).
Figure Extraction and Segmentation is the first step. We utilized Uni-Parser Fang2025, a specialized layout analysis tool for scientific documents, to process the source PDFs. Reaction schemes were automatically detected, cropped, and saved as high-resolution single images. Crucially, this process was designed to identify and include the scheme's associated titles and labels (subscripts) within the cropped area, thereby ensuring the completeness of the visual information. This approach guarantees that the visual inputs for the benchmark contain all necessary internal context while being isolated from the surrounding main text, focusing the task strictly on visual chemical perception and preventing information leakage from the body text.
The second step involves Hybrid Question Generation, where we adopted a “Human-in-the-Loop” strategy to produce high-quality Question-Answer (QA) pairs. Initially, OpenAI o4-mini Hurst2024 was seeded with specific chemical templates to propose a set of five candidate QA pairs per image. (The prompt was documented in Supporting Information S4). Domain experts then acted as “exam setters”, rigorously reviewing, modifying, or completely rewriting these questions to ensure chemical relevance. These questions were strictly categorized into two tiers to assess different levels of cognitive capability. The first tier, Descriptive Questions, focuses on information extraction, assessing the model's ability to accurately “read” the diagram. This encompasses tasks such as identifying reaction titles, enumerating reactants and products, and extracting precise reaction parameters like temperature, reaction time, and specific activation methods (e.g., microwave irradiation, ball milling, or photoredox conditions). It also involves distinguishing the roles of various species, such as differentiating a metal catalyst from its ligand. The second tier, Reasoning Questions, requires synthesizing visual information with chemical knowledge to derive conclusions. This ranges from simple reasoning, such as comparing yields and enantioselectivity across substrates or performing Markush enumeration, to complex reasoning tasks like mechanism analysis, identifying bond cleavage sites, and tracing catalytic cycles. Advanced inquiries further test the model's ability to analyze structure-activity relationships (SAR) based on electronic effects and perform retrosynthetic planning.
Following the validation of correct answers, we proceeded to construct a rigorous multiple-choice evaluation framework through a two-stage process involving automated generation and expert refinement. Initially, OpenAI o4-mini was utilized to propose three distractor options for each validated QA pair. However, recognizing that raw LLM-generated distractors are often plagued by hallucinations or formatting artifacts that make them easily identifiable, our domain experts performed a final round of “Adversarial Editing” to create “Hard Negatives”. This step was critical to ensuring that the benchmark tests genuine chemical understanding rather than the model's ability to exploit test-taking heuristics.
The adversarial refinement process adhered to the following principles. First, experts minimized hallucinations, removing chemically impossible valencies or non-existent functional groups to ensure all options represented valid chemical entities. To prevent “gaming” via pattern matching, we standardized the format and length of all options, minimizing “length bias” where correct answers might otherwise stand out due to excessive detail. Crucially, the distractors were designed to enforce visual dependency; options were constructed to be chemically plausible yet mutually exclusive based on the visual evidence (e.g., offering both E- and Z- isomers), compelling the model to analyze the specific reaction scheme rather than relying on general chemical knowledge. Finally, we maximized chemical plausibility by mimicking common human errors—such as including enantiomers, diastereomers, or regioisomers—thereby ensuring that high performance on RxnBench reflects deep chemical insight.
Construction of Full-Document QA (FD-QA)
The Full-Document QA (FD-QA) task represents the pinnacle of the RxnBench evaluation hierarchy, designed to test a model's ability to synthesize information from complete scientific articles—mimicking the daily workflow of a research chemist(Figure 2b). Based on 108 representative articles, we constructed a dataset that demands rigorous cross-modal reasoning rather than simple keyword matching. Unlike the single-figure task, FD-QA is formulated as a variable-response multiple-choice problem where each question offers four substantive options (A-D) containing zero to four correct answers. Crucially, to penalize models that rely on “hallucinated guessing”, we introduced a “None of the Above” option (Option E), which must be selected if no correct answer is present among the provided choices. This design forces the model to perform verification rather than mere selection.
A distinguishing feature of FD-QA, compared to the predominantly text-based options in SF-QA, is the diversity of option formats designed to comprehensively evaluate visual-linguistic alignment. Options are presented in three distinct modalities: pure text containing standard chemical descriptions or numerical values; pure structure images (e.g., cropped from ChemDraw) that require the model to directly perceive and compare visual graphs without intermediate textual translation; and mixed text-image layouts where text descriptions are interleaved with structural snippets, simulating the complex layout often found in reaction scope tables.
Domain experts were instructed to design five diverse questions per article, adhering strictly to the principle of “Context-Dependent Reasoning”. Rather than simply requiring the presence of multiple modalities, it ensures that answering necessitates synthesizing information distributed across the document's full context—bridging textual narratives, graphical schemes, and tabular data. These questions encompass four advanced capability dimensions: multimodal information extraction, which involves retrieving and structuring data like product lists and yields embedded within complex layouts; contextual summarization, which requires synthesizing scattered details to reconstruct experimental procedures; chemical logic and reasoning, for tasks demanding scientific deduction such as explaining regioselectivity based on a mechanistic discussion; and reference resolution, which tests the ability to navigate the document to map specific textual entity mentions to their corresponding structural representations in the schemes.
To ensure rigor, the distractors in FD-QA are designed as contextually relevant but factually incorrect “hard negatives.” Annotators ensured that all incorrect options were derived directly from the PDF content (e.g., using conditions from a different entry in the same table) rather than being fabricated, thereby preventing models from ruling out options simply based on unfamiliarity. Furthermore, to rigorously evaluate fine-grained structural reasoning, we introduced a specialized category of Image-based Distractors. These options consist of chemically valid structures generated via professional drawing software that are visually “plausible but incorrect.” They are intentionally crafted to bear high similarity to the ground truth while differing in subtle, easily overlooked details—such as the regio-position of substituents, variations in carbon chain length, or specific stereochemical configurations—forcing the model to perform precise graph-level structural verification rather than relying on coarse visual features.
Results and discussion
Task Categorization and Distribution in SF-QA
To provide a granular analysis of model capabilities, we developed a comprehensive taxonomy for the Single-Figure QA (SF-QA) dataset, classifying the 1,525 curated questions into six distinct types based on the cognitive skills required. The definitions and distribution of these task types are summarized in Figure 3.
Figure 3: Distribution of Question Types in the SF-QA Dataset. The dataset covers 6 balanced types of tasks, ranging from fundamental information extraction to advanced mechanistic reasoning and structural recognition.
Fact Extraction (37.2%, 568 Questions): This foundational category tests the model's ability to directly retrieve explicit textual or numerical information embedded within reaction schemes. Tasks include identifying reaction titles, extracting yield values, or reading reaction times and temperatures without requiring chemical interpretation. (e.g., “What is the title of this chemical reaction scheme?”)
Reagent Roles & Functions (16.5%, 252 Questions): Moving beyond OCR, this type requires the model to understand the chemical semantics of the text labels. Models must distinguish between reagents, catalysts, solvents, and additives, and correctly identify their functional roles (e.g., “Which molecule acts as the oxidant in this transformation?”).
Mechanism & Process (19.4%, 296 Questions): This category assesses the model’s grasp of the dynamic nature of chemical reactions. Questions focus on interpreting the reaction progression, including identifying key intermediates, tracing catalytic cycles, and understanding the sequence of mechanistic steps depicted in the scheme. (e.g., “Which of the following statements correctly describes the role of photocatalyst (PC) in the catalysis cycle?”)
Comparative Analysis & Reasoning (14.2%, 216 Questions): These questions require the model to perform comparative evaluations or causal reasoning. Typical tasks include predicting how changes in substrates (e.g., electronic effects of substituents) influence reaction outcomes like yield or selectivity, often necessitating an analysis of provided scope tables. (e.g., “Compare the effects of different LED power and wavelength on the reaction yield.”)
Global Understanding (3.9%, 59 Questions): Representing a higher-level challenge, this type tests the comprehension of multi-step synthetic pathways. Models must demonstrate an understanding of step-to-step coherence and the overall synthetic design strategy. (e.g., “Please describe the overall transformation type and the key reaction processes of the electrochemical reaction shown in the figure.”)
Structure Recognition (8.8%, 134 Questions): This task focuses on the precise translation of visual molecular graphs into machine-readable strings (SMILES or E-SMILES). Unlike general image captioning, this requires rigorous graph-level parsing to capture exact connectivity, stereochemistry, and functional group identity, serving as a direct evaluation of the model's “chemical vision.” (e.g., “What is the E-SMILES of the product containing a tert-butyl group?”)
As illustrated in Figure 3, while Fact Extraction constitutes the largest portion (37.2%) to ensure a solid baseline for visual perception, the majority of the benchmark (62.8%) is dedicated to tasks requiring varying degrees of chemical knowledge and reasoning. Notably, nearly one-fifth of the questions are dedicated to mechanism understanding, highlighting RxnBench's focus on deep chemical logic rather than superficial pattern matching.
Overall Model Performance on SF-QA
We evaluated a comprehensive suite of Multimodal Large Language Models (MLLMs), ranging from proprietary giants to state-of-the-art open-weight models, using a Zero-shot prompting strategy. To assess linguistic robustness, all models were queried using both English (RxnBench-En) and Chinese (RxnBench-Zh) prompts. To handle non-standard outputs where models typically generate verbose explanations instead of a specific option, we employed a GPT-4o-based mapping strategy to align free-text responses to the standard A–D options, ensuring a fair comparison across instruction-following capabilities.
Table 1: Performance comparison of selected representative MLLMs on the SF-QA dataset.
Models are ranked by Mean Score.
Model Thinka Weightb RxnBench-En RxnBench-Zh Mean Score Gemini-3-Flash-preview - 0.9593 0.9652 0.9623 Seed1.8-Think - 0.9325 0.9403 0.9364 Gemini-3-Pro-preview - 0.9318 0.9403 0.9361 GPT-5(high) - 0.9279 0.9246 0.9263 Gemini-2.5-Pro - 0.9095 0.9423 0.9259 Qwen3-VL-235B-A22B-Think Open 0.9220 0.9134 0.9177 Seed1.5-VL-Think - 0.9056 0.9161 0.9109 InternVL3.5-241B-A28B Open 0.9003 0.9062 0.9033 Intern-S1 Open 0.8938 0.8944 0.8941 Seed1.5-VL - 0.8518 0.8669 0.8594 Qwen3-VL-235B-A22B-Instruct Open 0.8492 0.8675 0.8584 GPT-4o - 0.7462 0.7436 0.7449 Random - - 0.2500 0.2500 0.2500
To comprehensively map the capability landscape, we evaluated a total of 41 Multimodal Large Language Models, spanning diverse architectures (from lightweight to massive-scale) and access types (proprietary vs. open-source). The overall accuracy results for selected representative models are summarized in Table 1 (see Supplementary Table S1 for the complete data of all evaluated models). Overall, the evaluation highlights a performance hierarchy driven by model scale and inference-time reasoning, where proprietary frontiers and “thinking” models lead, yet open-source alternatives are rapidly narrowing the gap with robust bilingual consistency.
As shown in Table 1, the leaderboard is currently dominated by next-generation proprietary models. Gemini-3-Flash-preview achieves the highest mean score of 96.23%, demonstrating exceptional capability in chemical vision understanding. It is closely followed by Seed1.8-Think (93.64%) and Gemini-3-Pro-preview (93.61%). These results suggest that the scaling laws of model size and data quality continue to drive significant improvements in scientific domain tasks.
A notable trend is the superior performance of models equipped with inference-time reasoning capabilities (denoted as “Think” in Table 1). For instance, Compare with Qwen3-VL-235B-A22B-Instruct (85.84%), Qwen3-VL-235B-A22B-Think achieves a mean score of 91.77%, effectively bridging the gap between open-weight and top-tier proprietary models. It indicates that allowing the model to generate internal chains of thought before outputting an answer is particularly beneficial for the complex, multi-step reasoning required in chemical reaction analysis.
While proprietary models lead the pack, open-source models have made remarkable strides. The Qwen3-VL and InternVL series perform competitively, with the best open models surpassing previous generation proprietary models like GPT-4o (74.49%) by a significant margin. This democratization of capability is promising for the deployment of local AI chemists. Furthermore, most top-tier models exhibit high consistency between English and Chinese benchmarks (e.g., Gemini-3-Flash-preview: 95.93% EN vs 96.52% ZH), reflecting strong cross-lingual alignment in their pre-training, though a slight performance drop is observed in some smaller models when processing Chinese prompts.
Detailed Analysis by Question Type on SF-QA
To dissect the fine-grained capabilities of current MLLMs, we decomposed the SF-QA performance across the six defined task types. The results, detailed in Table 2(see Supplementary Table S2 for the complete data of all evaluated models), reveal significant variance in model proficiency depending on the cognitive depth required.
A distinct capability gap emerges as tasks increase in complexity. For fundamental Fact Extraction, top-tier models demonstrate near-perfect proficiency, with Gemini-3-Pro-preview achieving 96.48% and Gemini-3-Flash-preview reaching 96.13%, indicating that basic information retrieval is effectively solved. However, in complex reasoning tasks, a clear divergence appears. Standard baselines like GPT-4o show limited capability (scoring 79.73% and 76.27% respectively), whereas models equipped with inference-time reasoning maintain exceptional robustness. For instance, GPT-5 (high) achieves a remarkable 96.61% on Global Understanding, and Gemini-3-Flash-preview achieves 97.64% on Mechanism & Process. This performance contrast highlights a “reasoning leap”, where internal chains of thought successfully bridge the gap between perception and deep chemical logic.
Despite the reasoning advantage, a critical performance drop reappears in Structure Recognition across the board. Even leading proprietary models struggle to match their text-based performance in this category; for example, Gemini-3-Pro-preview drops to 74.63%, and GPT-4o falls to 52.24%. Although specific architectures like Qwen3-VL-235B-A22B-Think demonstrate a relative advantage with 84.33% and Gemini-3-Flash-preview reaches 90.30%, the general downward trend across most models suggests that reasoning alone cannot fully compensate for the deficits in visual molecular encoding for all architectures (Figure 4).
Collectively, these findings indicate that while current SOTA models have mastered the textual “reading” of chemical literature, the frontiers of accurate visual molecular perception and deep mechanistic “understanding” remain active challenges. The data suggests that future advancements must prioritize domain-specific visual encoders to resolve the universal bottleneck in structure recognition.
Table 2: Detailed Model Performance selected across SF-QA Task Types
Model Thinka Weightb F.E.c RR&F.d M&P.e CA&R.f G.U.g S.R.h
Gemini-3-Flash-preview - 0.9613 0.9643 0.9764 0.9630 0.9492 0.9030
Seed1.8-Think - 0.9331 0.9484 0.9527 0.9444 0.9492 0.8284
Gemini-3-Pro-preview - 0.9648 0.9246 0.9527 0.9398 0.9322 0.7463
GPT-5(high) - 0.9313 0.9444 0.9527 0.9167 0.9661 0.8358
Gemini-2.5-Pro - 0.9331 0.9246 0.9459 0.9491 0.9322 0.6343
Qwen3-VL-235B-A22B-Think Open 0.9190 0.9405 0.9459 0.9213 0.9322 0.8433
Seed1.5-VL-Think - 0.8996 0.9365 0.9358 0.9074 0.9153 0.8060
InternVL3.5-241B-A28B Open 0.8944 0.9127 0.9291 0.9167 0.9153 0.8134
Intern-S1 Open 0.9014 0.9127 0.9223 0.9028 0.8814 0.7463
Seed1.5-VL - 0.9327 0.9127 0.9122 0.8472 0.8305 0.7015
Qwen3-VL-235B-A22B-Instruct Open 0.8204 0.8929 0.8986 0.8426 0.8814 0.7761
GPT-4o - 0.7359 0.8175 0.7973 0.7500 0.7627 0.5224
Random - - 0.2500 0.2500 0.2500 0.2500 0.2500 0.2500
  • a
    Think: indicates models with inference-time reasoning capabilities
  • b
    Weight: distinguishes between Proprietary (Closed) and Open-weight models
  • c
    F.E.: Fact Extraction
  • d
    RR&F.: Reagent Roles & Functions
  • e
    M&P.: Mechanism & Process
  • f
    CA&R.: Comparative Analysis & Reasoning
  • g
    G.U.: Global Understanding
  • h
    S.R.: Structure Recognition
Table 3: Detailed Model Performance selected across FD-QA Task Types
Model Thinka Weightb Rxn.En.c Rxn.Zh.d C.R.e S.R.f O.S.g
Gemini-3-Flash-preview - 0.4574 0.4685 0.5530 0.4264 0.4630
Gemini-3-Pro-preview - 0.4167 0.4407 0.5000 0.3897 0.4287
Gemini-2.5-Pro - 0.4148 0.4426 0.5682 0.3652 0.4287
Gemini-2.5-Flash - 0.3519 0.3815 0.4394 0.3235 0.3667
Seed1.8-Think - 0.3352 0.3889 0.4242 0.3064 0.3620
GPT-5(high) - 0.3444 0.3463 0.4470 0.3113 0.3454
Seed1.5-VL-Think - 0.3148 0.3574 0.3939 0.2892 0.3361
Qwen3-VL-235B-A22B-Think Open 0.3056 0.3056 0.4091 0.2721 0.3056
Intern-S1 Open 0.1963 0.1944 0.2652 0.1740 0.1954
Seed1.5-VL-Instruct - 0.1630 0.1741 0.2045 0.1495 0.1685
GPT-4o - 0.1426 0.1389 0.1515 0.1397 0.1407
Qwen3-VL-235B-A22B-Instruct Open 0.1241 0.1481 0.1742 0.1078 0.1361
  • a
    Think: indicates models with inference-time reasoning capabilities
  • b
    Weight: distinguishes between Proprietary (Closed) and Open-weight models
  • c
    Rxn.En.: RxnBench-Doc-En
  • d
    Rxn.Zh.: RxnBench-Doc-Zh
  • e
    C.R.: Context Reasoning (En)
  • f
    S.R.: Structure Reasoning (En)
  • g
    O.S.: Overall Score
Evaluation on Full-Document QA (FD-QA)
Originated from 108 research articles, the FD-QA benchmark comprises 540 meticulously curated questions designed to assess document-level chemical comprehension. These tasks are categorized into two distinct types: Context Reasoning (24.4%), which requires the integration of information from multimodal sources such as reaction images, tables, and text to answer questions; and Structure Reasoning (75.6%), which focuses on rigorous logical deductions related to molecular and Markush structures, reaction components, and mechanistic inference.
Figure 4: Comparative performance analysis of inference-time reasoning (“Think”) versus standard instruction-tuned (“Instruct”) models across diverse cognitive tasks. The visualization highlights the “reasoning leap” in complex mechanistic and synthetic tasks achieved by thinking models (red/green lines) compared to standard baselines (blue/yellow lines), while also illustrating the universal bottleneck in precise structure recognition.
To simulate a realistic reading environment, we render each PDF document into a sequence of high-resolution images at 144 dpi, serving as inputs to the MLLM. The <answer> tag in the question acts as a placeholder that is replaced by the corresponding image list, forming an interleaved vision–language prompt. We then employ a GPT-4o-based extraction pipeline to parse the model’s raw output into a standard multiple-select format (A–E), explicitly mapping statements such as “None of the above” to Option E. The extracted answers are compared against the ground truth using an exact-match criterion, where any missing or extraneous selections are counted as errors. We report absolute accuracy for both the English and Chinese versions. Note that cases where the model exceeds the allowable context length at inference time result in missing scores, an issue observed in part of the Intern-S1 results.
We evaluated a total of 23 Multimodal Large Language Models. The results, summarized in Table 3, reveal that FD-QA represents a significantly higher difficulty tier compared to single-figure tasks. Note that due to space constraints, Table 3 highlights the performance of selected representative models that define the current state-of-the-art or serve as key baselines; the complete leaderboard featuring all 23 evaluated models is provided in Tabel S3 in Supplementary Information.
The evaluation exposes a stark performance chasm where models lacking inference-time reasoning fail to surpass a 15% accuracy threshold, whereas “thinking” models (particularly the Gemini family which leads with ~46% accuracy) demonstrate superior capability in handling long-context multimodal inputs, effectively validating the necessity of Chain-of-Thought processing for complex document synthesis. Even so, the fact that no model reaches 50% accuracy underscores the substantial challenge of synthesizing full scientific narratives compared to isolated perception tasks.
Crucially, a significant disparity exists between reasoning modalities. As shown in the detailed breakdown, models consistently underperform on Structure Reasoning compared to Context Reasoning (e.g., Gemini-2.5-Pro: 36.52% vs 56.82%). This trend persists across almost all evaluated models (Figure 5), indicating that while MLLMs are relatively adept at retrieving and correlating textual or tabular information (context), their ability to perform precise logical deductions based on visual molecular structures remains a distinct bottleneck.
Figure 5: Comprehensive performance evaluation of MLLMs on FD-QA benchmark. The chart details the accuracy of various models across Context Reasoning and Structure Reasoning. The results highlight the substantial advantage of “thinking” models over standard architectures in handling long-context scientific synthesis, while also revealing a consistent performance lag in structure-dependent reasoning tasks.
Conclusion
In this work, we introduced RxnBench, a pioneering multimodal benchmark meticulously designed to evaluate the chemical reaction understanding capabilities of Large Language Models within the context of authentic scientific literature. By constructing a dual-tiered evaluation framework comprising Single-Figure QA (SF-QA) and Full-Document QA (FD-QA), we have provided the community with a rigorous standard to assess not just perceptual accuracy, but the depth of chemical reasoning.
Our comprehensive evaluation reveals a distinct landscape where MLLMs have mastered surface-level information extraction but face substantial bottlenecks in deep chemical reasoning and precise structural recognition—critical deficits that are partially mitigated by inference-time reasoning capabilities yet remain acute when synthesizing complex, long-context scientific narratives.
RxnBench serves as a diagnostic tool that points the way forward. Future research should prioritize the development of domain-specific visual encoders capable of pixel-perfect molecular graph recognition and the integration of external chemical tools (e.g., RDKit, calculator agents) to verify hallucinations. Furthermore, moving beyond passive QA to active Agentic Workflows—where models autonomously navigate, query, and verify literature data—will be essential for realizing the vision of a truly autonomous AI Chemist.
Associated content
Data Availability Statement
SF-QA Benchmark dataset is available at https://huggingface.co/datasets/UniParser/RxnBench
FD-QA Benchmark dataset is available at https://huggingface.co/datasets/UniParser/RxnBench-Doc.
Acknowledgements
This research was supported by the New Generation Artificial Intelligence-National Science and Technology Major Project 2025ZD0121905.
We thank the students from Shanghai Jiao Tong University who contributed to the establishment of RxnBench (No particular order): Jing Guan, Jingli He, Panpan Li, Yifan Song, Tiandong Guan, Guanlin Li, Jiacheng Zhang, Qinyi Li, Shaoyi Dan, Lefeng Dong, Xintai Zhang, Xuezhen Kou, Jie Zhao, Yicong Luo, and Haoyang Wang.
Table S1: Performance comparison of representative MLLMs on the SF-QA.
Model Think Weight RxnBench-En RxnBench-Zh Mean Score
Gemini-3-Flash-preview - 0.9593 0.9652 0.9623
Seed1.8-Think - 0.9325 0.9403 0.9364
Gemini-3-Pro-preview - 0.9318 0.9403 0.9361
GPT-5(high) - 0.9279 0.9246 0.9263
Gemini-2.5-Pro - 0.9095 0.9423 0.9259
GPT-5.1(high) - 0.9213 0.9220 0.9216
GPT-5(medium) - 0.9207 0.9226 0.9216
Qwen3-VL-235B-A22B-Think Open 0.9220 0.9134 0.9177
Qwen3-VL-32B-Think Open 0.9128 0.9161 0.9144
GPT-5.1(medium) - 0.9108 0.9141 0.9125
GPT-5-mini - 0.9108 0.9128 0.9118
Seed1.5-VL-Think - 0.9056 0.9161 0.9109
o3 - 0.9056 0.9115 0.9086
o4-mini - 0.9062 0.9075 0.9069
InternVL3.5-241B-A28B Open 0.9003 0.9062 0.9033
Intern-S1 Open 0.8938 0.8944 0.8941
Qwen3-VL-30B-A3B-Think Open 0.8689 0.8590 0.8689
Qwen3-VL-Plus - 0.8551 0.8656 0.8604
Qwen3-VL-8B-Think Open 0.8636 0.8564 0.8600
Seed1.5-VL - 0.8518 0.8669 0.8594
Qwen3-VL-235B-A22B-Instruct Open 0.8492 0.8675 0.8584
InternVL3-78b Open 0.8531 0.8308 0.8420
Qwen3-VL-4B-Think Open 0.8577 0.8256 0.8416
Intern-S1-mini Open 0.8521 0.8282 0.8402
GLM-4.1V-9B-Thinking Open 0.8392 0.8341 0.8367
Qwen3-VL-32B-Instruct Open 0.8315 0.8407 0.8361
Qwen2.5-VL-72B Open 0.8341 0.8308 0.8325
Qwen2.5-VL-Max - 0.8192 0.8262 0.8227
GPT-5-nano - 0.7980 0.7941 0.7961
Qwen2.5-VL-32B Open 0.7980 0.7908 0.7944
Gemini-2.5-Flash - 0.6925 0.8557 0.7741
Qwen3-VL-8B-Instruct Open 0.7548 0.7495 0.7521
Qwen3-VL-30B-A3B-Instruct Open 0.7456 0.7436 0.7456
GPT-4o - 0.7462 0.7436 0.7449
Qwen2.5-VL-7B Open 0.7082 0.7233 0.7158
Qwen3-VL-4B-Instruct Open 0.7023 0.7023 0.7023
Qwen3-VL-2B-Think Open 0.6780 0.6708 0.6744
Qwen2.5-VL-3B Open 0.6748 0.6643 0.6696
GPT-4o mini - 0.6636 0.6066 0.6351
Qwen3-VL-2B-Instruct Open 0.5711 0.5928 0.5820
Choose the longest answer - - 0.4262 0.4525 0.4394
Deepseek-VL2 Open 0.4426 0.4216 0.4321
Random - - 0.2500 0.2500 0.2500
Table S2: Performance of evaluated models across SF-QA Task Types
Model Thinka Weightb F.E.c RR&F.d M&P.e CA&R.f G.U.g S.R.h
Gemini-3-Flash-preview - 0.9613 0.9643 0.9764 0.9630 0.9492 0.9030
Seed1.8-Think - 0.9331 0.9484 0.9527 0.9444 0.9492 0.8284
Gemini-3-Pro-preview - 0.9648 0.9246 0.9527 0.9398 0.9322 0.7463
GPT-5(high) - 0.9313 0.9444 0.9527 0.9167 0.9661 0.8358
Gemini-2.5-Pro - 0.9331 0.9246 0.9459 0.9491 0.9322 0.6343
GPT-5.1(high) - 0.9243 0.9524 0.9426 0.9167 0.9661 0.7910
GPT-5(medium) - 0.9349 0.9325 0.9493 0.9167 0.9492 0.7761
Qwen3-VL-235B-A22B-Think Open 0.9190 0.9405 0.9459 0.9213 0.9322 0.8433
Qwen3-VL-32B-Think Open 0.9296 0.9405 0.9426 0.9259 0.9153 0.7015
GPT-5.1(medium) - 0.9243 0.9365 0.9426 0.9167 0.9492 0.7090
GPT-5-mini - 0.9225 0.9325 0.9257 0.9259 0.9831 0.7388
Seed1.5-VL-Think - 0.8996 0.9365 0.9358 0.9074 0.9153 0.8060
o3 - 0.9313 0.9325 0.9223 0.8981 0.9492 0.7090
o4-mini - 0.6391 0.7302 0.7500 0.6667 0.6271 0.4627
InternVL3.5-241B-A28B Open 0.8944 0.9127 0.9291 0.9167 0.9153 0.8134
Intern-S1 Open 0.9014 0.9127 0.9223 0.9028 0.8814 0.7463
Qwen3-VL-30B-A3B-Think Open 0.8732 0.8810 0.9054 0.8843 0.9322 0.6940
Qwen3-VL-Plus - 0.8275 0.8968 0.8986 0.8565 0.9153 0.7687
Qwen3-VL-8B-Think Open 0.8768 0.8730 0.8885 0.9028 0.8983 0.6567
Seed1.5-VL - 0.9327 0.9127 0.9122 0.8472 0.8305 0.7015
Qwen3-VL-235B-A22B-Instruct Open 0.8204 0.8929 0.8986 0.8426 0.8814 0.7761
InternVL3-78b Open 0.8556 0.8730 0.8885 0.8981 0.9153 0.6194
Qwen3-VL-4B-Think Open 0.8838 0.8770 0.8615 0.9074 0.8983 0.6045
Intern-S1-mini Open 0.8239 0.8690 0.8547 0.8611 0.8475 0.6791
GLM-4.1V-9B-Thinking Open 0.8433 0.8690 0.8649 0.8657 0.8814 0.6493
Qwen3-VL-32B-Instruct Open 0.8169 0.8571 0.8885 0.8519 0.8305 0.6866
Qwen2.5-VL-72B Open 0.8063 0.8063 0.8770 0.9088 0.8102 0.9322
Qwen2.5-VL-Max - 0.7958 0.8571 0.8885 0.8194 0.8983 0.6642
GPT-5-nano - 0.8063 0.8452 0.8311 0.8241 0.7797 0.5672
Qwen2.5-VL-32B Open 0.7729 0.8413 0.8750 0.8009 0.8305 0.6418
Gemini-2.5-Flash - 0.7799 0.6111 0.6757 0.6620 0.7627 0.5373
Qwen3-VL-8B-Instruct Open 0.7113 0.8175 0.8446 0.8241 0.7627 0.5075
Qwen3-VL-30B-A3B-Instruct Open 0.7042 0.7937 0.8311 0.7824 0.7119 0.5970
GPT-4o - 0.7359 0.8175 0.7973 0.7500 0.7627 0.5224
Qwen2.5-VL-7B Open 0.6678 0.7659 0.8041 0.7130 0.6441 0.5373
Qwen3-VL-4B-Instruct Open 0.6708 0.7302 0.7804 0.7222 0.6610 0.5970
Qwen3-VL-2B-Think Open 0.7342 0.6706 0.7128 0.7083 0.6102 0.3657
Qwen2.5-VL-3B Open 0.6426 0.7381 0.7635 0.6898 0.6610 0.4776
GPT-4o mini - 0.6391 0.7302 0.7500 0.6667 0.6271 0.4627
Qwen3-VL-2B-Instruct Open 0.5405 0.6190 0.6318 0.6250 0.6102 0.3731
Deepseek-VL2 Open 0.4120 0.5040 0.4899 0.4907 0.3729 0.3060
Random - - 0.2500 0.2500 0.2500 0.2500 0.2500 0.2500
  • a
    Think: indicates models with inference-time reasoning capabilities
  • b
    Weight: distinguishes between Proprietary (Closed) and Open-weight models
  • c
    F.E.: Fact Extraction
  • d
    RR&F.: Reagent Roles & Functions
  • e
    M&P.: Mechanism & Process
  • f
    CA&R.: Comparative Analysis & Reasoning
  • g
    G.U.: Global Understanding
  • h
    S.R.: Structure Recognition
Table S3: Performance of MLLMs across FD-QA Task Types
Model Thinka Weightb Rxn.En.c Rxn.Zh.d C.R.e S.R.f O.S.g
Gemini-3-Flash-preview - 0.4574 0.4685 0.5530 0.4264 0.4630
Gemini-3-Pro-preview - 0.4167 0.4407 0.5000 0.3897 0.4287
Gemini-2.5-Pro - 0.4148 0.4426 0.5682 0.3652 0.4287
Gemini-2.5-Flash - 0.3519 0.3815 0.4394 0.3235 0.3667
Seed1.8-Think - 0.3352 0.3889 0.4242 0.3064 0.3620
GPT-5(high) - 0.3444 0.3463 0.4470 0.3113 0.3454
GPT-5.1(high) - 0.3333 0.3426 0.4545 0.2941 0.3380
Seed1.5-VL-Think - 0.3148 0.3574 0.3939 0.2892 0.3361
Qwen3-VL-32B-Think Open 0.2981 0.3167 0.4015 0.2647 0.3074
Qwen3-VL-235B-A22B-Think Open 0.3056 0.3056 0.4091 0.2721 0.3056
Qwen3-VL-30B-A3B-Think Open 0.2611 0.2481 0.3409 0.2353 0.2546
Qwen3-VL-8B-Think Open 0.2333 0.2315 0.3333 0.2010 0.2324
Qwen3-VL-4B-Think Open 0.2352 0.1870 0.3106 0.2108 0.2111
Intern-S1 Open 0.1963 0.1944 0.2652 0.1740 0.1954
Seed1.5-VL-Instruct x - 0.1630 0.1741 0.2045 0.1495 0.1685
Qwen3-VL-2B-Think Open 0.1444 0.1463 0.1818 0.1324 0.1453
GPT-4o - 0.1426 0.1389 0.1515 0.1397 0.1407
Qwen3-VL-235B-A22B-Instruct Open 0.1241 0.1481 0.1742 0.1078 0.1361
Qwen3-VL-32B-Instruct Open 0.1241 0.0981 0.1894 0.1029 0.1111
Qwen3-VL-8B-Instruct Open 0.0907 0.1019 0.1591 0.0686 0.0963
Qwen3-VL-30B-A3B-Instruct Open 0.0778 0.0815 0.1288 0.0613 0.0796
Qwen3-VL-4B-Instruct Open 0.0611 0.0611 0.0606 0.0612 0.0611
Qwen3-VL-2B-Instruct Open 0.0481 0.0667 0.0758 0.0392 0.0574
  • a
    Think: indicates models with inference-time reasoning capabilities
  • b
    Weight: distinguishes between Proprietary (Closed) and Open-weight models
  • c
    Rxn.En.: RxnBench-Doc-En
  • d
    Rxn.Zh.: RxnBench-Doc-Zh
  • e
    C.R.: Context Reasoning (En)
  • f
    S.R.: Structure Reasoning (En)
  • g
    O.S.: Overall Score
 S4: Prompt used in SF-QA generation
Role@ :
Chemistry Expert (Organic \& Cheminformatics).
bTask@ :
Analyze a reaction scheme image and generate one high-difficulty, reasoning-based MCQ in JSON format.
Requirements@ :
Content: Question must be grounded strictly in the image (SMILES for concrete, Extended-SMILES for Markush). Focus on mechanism, multi-step logic, or structural inference.
Options: 1 correct answer, 3 plausible but incorrect distractors. All must match in tone, length, and style.
Language: Provide both Chinese and English for all text fields.
Taxonomy: Select exactly one: 0 (Fact Extraction), 1 (Reagent Roles and Functions Identification), 2 (Reaction Mechanism and Process Understanding), 3 (Comparative Analysis and Reasoning), 4 (Multi-step Synthesis and Global Understanding), 5 (Chemical Structure Recognition).
JSON Output Only:
‘‘‘ json
"type": "X", \# from 0 to 5
"question\_zh": "question in Chinese",
"answer\_zh": "correct answer in Chinese",
"wrong\_options\_zh": {[}"wrong option 1 in Chinese", "wrong option 2 in
Chinese", "wrong option 3 in Chinese"{]},
"question\_en": "question in English",
"answer\_en": "correct answer in English",
"wrong\_options\_en": {[}"wrong option 1 in English", "wrong option 2 in
English", "wrong option 3 in English"{]}
bConstraints@ :
- No external knowledge.
- Single unambiguous answer.
- No markdown/commentary outside JSON.
- Do not use phrases like "as shown in the image".